# Forging GEMs: Advancing Greek NLP through Quality-Based Corpus Curation and Specialized Pre-training

### v1

@@ Abstract
The advancement of natural language processing for morphologically rich, moderately-resourced languages like Modern Greek is often hindered by a fragmented research landscape, a lack of architectural diversity and reliance on limited context-length models. This is particularly true in specialized, high-value domains such as law, where existing models are frequently confined to early transformer architectures with a restrictive 512-token window, insufficient for analyzing long legal documents. To address these challenges, this paper presents Greek Embedding Models, a new family of transformer models for Greek language built upon a foundation of extensive, quality-driven data curation. We detail the construction of several large-scale Greek corpora, emphasizing a rigorous, quality-based filtering and preprocessing methodology to create high-value training datasets from both general-domain and specialized legal sources. On this carefully curated foundation, we pre-train and systematically evaluate a diverse suite of modern architectures, which has not previously applied to Greek language, such as ELECTRA, ConvBERT and ModernBERT. Furthermore, we propose the first bilingual Greek-English Embedding Models tailored for the legal domain. The extensive experiments on downstream tasks demonstrate that the new class of models establish the effectiveness of the proposed approach, highlighting that the GEM-RoBERTa and GEM-ConvBERT models significantly outperform existing baselines.
@@ 1 Introduction
The advent of transformer-based language models [43, 12] has revolutionized Natural Language Processing (NLP), establishing a dominant paradigm of large-scale pre-training followed by task-specific fine-tuning. While foundational models were initially developed for English, subsequent research has consistently shown that performance is significantly enhanced by creating monolingual models for other languages and adapting them to high-value domains, such as law (e.g., for Greek [31, 4] and German [11]) or biomedicine (e.g., for French [25]). For moderately-resourced languages such as Greek, this specialization is crucial, as dedicated monolingual models like Greek-BERT [23] have proven superior to their multilingual counterparts on a range of NLP tasks.
@@
The Greek NLP ecosystem has since evolved rapidly along several parallel tracks. Following the foundational Greek-BERT [23], a line of domain-specific models for the legal field emerged, including the two versions of Greek Legal BERT [31, 4] and the architecturally superior Greek-Legal RoBERTa [38], which progressively advanced the state-of-the-art on Greek legal tasks. Concurrently, separate efforts addressed this limitation for general-domain Greek with models like the Greek Longformer, which supports a 4096-token context window [51] and Meltemi-7B [45], a large-scale generative model with an 8192-token context. However, the divergent progress, has created a fragmented landscape where the benefits of modern architectures and extended context lengths have not been systematically unified for specialized domains like Greek law.
@@
Despite this progress, further advancement in Greek NLP is hindered by several fundamental limitations. Firstly, the development of high-quality, large-scale datasets remains a significant challenge, often leading to sub-optimal model performance. Secondly, researchers also frequently neglect the inherent linguistic complexity of morphologically rich languages like Greek, where intricate syntactic structures and domain-specific terminology demand special attention. This leads to a third limitation: a lack of architectural innovation. The Greek NLP space and the legal domain in particular, has been confined to early encoder architectures like BERT and RoBERTa, often with a restrictive 512-token context length, while more efficient and powerful models such as ELECTRA [9] and DeBERTa [15] have not been explored. This architectural stagnation prevents the community from leveraging state-of-the-art advancements effectively.
@@
This work is motivated by the need to systematically address these challenges. We bridge the existing gaps by undertaking the first comprehensive pre-training and evaluation of a diverse suite of modern transformer architectures for the Greek language, with a specific focus on the complex legal domain as a proving ground. We specifically target this domain as its highly structured, terminologically rich and stylistically consistent language serves as a low-noise, high-signal corpus, ideal for effectively capturing the complex grammar and syntax of Greek. We propose a new family of Greek Embedding Models (GEMs), pre-trained from scratch on several meticulously curated corpora, including both a large-scale general-purpose dataset and specialized domain-specific ones. Our approach explores architectural innovations not previously applied to Greek and is the first to systematically train and evaluate a suite of Greek models with an extended 1024-token context window. Furthermore, we investigate data-centric training strategies to enhance model performance.
In summary, the main contributions of this work are:
@@
  • We investigate data-centric strategies, showing that targeted repetition of high-quality legal sub-corpora is an effective strategy for domain adaptation. We also introduce the first bilingual Greek-English Embedding Models (GEM) specifically tailored for the legal domain.
@@
The remainder of this work is organized as follows: Section 2 reviews related work in Greek language modeling. Section 3 describes the composition of the datasets. Section 4 details the experimental setup, including the model architectures, corpora configurations and training procedures. Section 5 presents the results of the numerical experiments and analysis. Finally, Section 6 discusses our findings and concludes the paper.
@@ 2 Related Work
The development of effective pre-trained language models for morphologically rich languages has emerged as a fundamental challenge in advancing natural language processing capabilities beyond English-centric paradigms [36, 35]. This challenging task has been primarily driven by two complementary research directions: (i) foundational architectural advancements in transformer-based models and their optimization strategies [43, 12, 27] and (ii) the systematic adaptation of these innovations to specific languages and specialized domains [30, 8, 5]. The first direction encompasses sophisticated developments in model architectures, training objectives and efficiency improvements, ranging from BERT’s masked language modeling [12] to more recent innovations such as ELECTRA’s replaced token detection [9] and DeBERTa’s disentangled attention mechanisms [15]. The second direction focuses on applying these architectural advances to non-English languages and domain-specific applications, where monolingual and specialized models consistently demonstrate superior performance over their multilingual counterparts [44, 20]. In the remainder of this section, we provide a comprehensive review of both areas, highlighting their contributions and limitations which motivate our proposed systematic approach to Greek language model development.
@@ 2.1 Architectural Innovations in Pre-trained Language Models
The evolution of transformer-based language models has been characterized by continuous architectural refinements aimed at improving training efficiency, model capability and contextual understanding. The foundational work began with the transformer architecture [43] and BERT’s [12] Masked Language Model (MLM) objective, which established the dominant paradigm of large-scale pre-training followed by task-specific fine-tuning. This seminal work demonstrated the effectiveness of bidirectional encoding and self-supervised learning for capturing contextual representations across diverse NLP tasks.
@@
Building upon BERT’s foundation, RoBERTa [27] introduced critical optimization improvements to the pre-training methodology, including the removal of the Next Sentence Prediction objective, dynamic masking strategies and extended training regimens. These modifications resulted in substantial performance gains across multiple benchmarks, establishing the importance of training procedure optimization alongside architectural innovation. The success of RoBERTa highlighted that careful attention to training dynamics could yield significant improvements without fundamental architectural changes.
@@
ELECTRA [9] introduced a paradigm shift through its sample-efficient Replaced Token Detection (RTD) objective, which trains the model to distinguish between original and replaced tokens rather than predicting masked tokens. This approach enables the model to learn from all input positions rather than only the masked subset, resulting in improved sample efficiency and computational effectiveness. The RTD objective demonstrated particular advantages in scenarios with limited computational resources while maintaining competitive performance across downstream tasks.
@@
DeBERTa [15] advanced the field through its disentangled attention mechanism, which separately encodes content and position information before combining them in attention computation. This architectural innovation, combined with enhanced mask decoder and absolute position embedding strategies, yielded substantial improvements in model performance while providing better interpretability of attention patterns. The disentangled approach proved particularly effective for tasks requiring fine-grained understanding of positional relationships and syntactic structures.
@@
The challenge of processing longer sequences was addressed by Longformer [6], which introduced sparse attention patterns to handle extended contexts efficiently. By combining local windowed attention, global attention for task-specific tokens and dilated attention patterns, Longformer successfully scaled transformer models to process documents with thousands of tokens while maintaining computational tractability. This innovation proved crucial for applications requiring document-level understanding and long-range dependency modeling.
@@
ConvBERT [19] introduced a novel hybrid approach which combines convolutional and self-attention mechanisms to improve efficiency and effectiveness. The model replaces some self-attention heads with cheaper convolutional operations, which capture local dependencies more efficiently while maintaining the global modeling capabilities of attention mechanisms. This architectural innovation reduces computational costs while often achieving superior performance, particularly on tasks requiring fine-grained local pattern recognition alongside global context understanding.
@@
Most recently, ModernBERT [47] has emerged as a comprehensive modernization of the original BERT architecture, incorporating lessons learned from years of research into transformer optimization. ModernBERT integrates several architectural improvements including GeGLU activation functions, RMSNorm normalization, unpadding techniques for efficiency and optimized attention implementations. The model demonstrates that systematic incorporation of incremental improvements can yield substantial performance gains while maintaining compatibility with existing BERT-based infrastructure, representing a mature synthesis of architectural innovations developed over the transformer era.
@@ 2.2 Greek Language Models and Domain Specialization
The development of Greek natural language processing capabilities has followed a progressive trajectory from foundational monolingual models to specialized domain-specific applications. Koutsikakis et al. [23] conducted the pioneering effort in modern Greek NLP by developing Greek BERT, the first monolingual BERT model specifically for the Greek language. Pre-trained on a comprehensive 29GB corpus comprising general-domain text from diverse sources, Greek BERT established crucial performance baselines and demonstrated superior results on tasks such as named entity recognition and natural language inference compared to multilingual models. This foundational work highlighted the significant benefits of language-specific pre-training for morphologically rich languages, where specialized tokenization and vocabulary development prove essential for capturing linguistic nuances.
@@
The success of Greek BERT catalyzed rapid development in domain-specific applications, particularly within the legal field. Papaloukas et al. [31] introduced the first Greek Legal BERT by adapting the original Greek BERT architecture through continued pre-training on a 5GB corpus of legal documents. This domain adaptation strategy demonstrated substantial improvements on legal-specific tasks, establishing the effectiveness of specialized pre-training for capturing domain-specific terminology and discourse patterns. The model’s superior performance on legal topic classification and named entity recognition tasks validated the importance of domain-focused training data curation.
@@
Apostolopoulou et al. [4] extended this work by developing an enhanced version of Greek Legal BERT which incorporated a larger and more diverse legal corpus, including additional sources of legal documents and regulatory texts. The expanded training data provided broader coverage of legal language variations and contemporary legal discourse, resulting in improved performance metrics across multiple legal NLP tasks. This work demonstrated that systematic corpus expansion and diversification constitute effective strategies for enhancing domain-specific model capabilities.
@@
Saketos et al. [38] represented the next significant advancement by moving beyond the original BERT architecture and leveraging RoBERTa’s optimized pre-training methodology for Greek legal applications. Greek-Legal RoBERTa demonstrated substantial performance gains over the Greek Legal BERT variants on identical downstream legal tasks, confirming that architectural improvements and training optimizations provide complementary benefits to specialized domain adaptation. The success of this approach established that both architectural innovation and domain-specific data curation are essential components for advancing specialized NLP capabilities.
@@
Parallel developments addressed the limitation of context length processing for general-domain Greek applications. The Greek Longformer [51] introduced extended context capabilities with a 4096-token window, enabling processing of longer Greek documents while maintaining computational efficiency through sparse attention mechanisms. Similarly, Meltemi-7B [45], a large-scale generative model, further extended context capabilities to 8192 tokens, demonstrating the feasibility of long-context modeling for Greek language applications. However, these general-domain advances remained isolated from specialized applications, creating a disconnect between architectural capabilities and domain-specific requirements.
@@ 2.3 Research Gaps and Contributions
The Greek NLP landscape is fragmented, characterized by several limitations that impede systematic progress. A key issue is the lack of a concerted effort to create comprehensive, large-scale and high-quality datasets that can serve both general-purpose and specialized applications. This absence of systematically curated training corpora fundamentally limits the potential for developing robust models. Secondly, the application of different model architectures has been piecemeal; while foundational models exist, more recent architectures such as ELECTRA, ModernBERT, DeBERTa and ConvBERT have not been systematically employed and evaluated for Greek language. Finally, a notable gap exists in context length processing for specialized fields. While general-domain models have achieved significantly longer context capabilities, specialized domains like legal NLP have yet to adopt these extensions, limiting their effectiveness on the long documents characteristic of legal and regulatory texts.
This work addresses these fundamental limitations by presenting Greek Embedding Models (GEMs), a new family of models developed through a comprehensive and systematic approach. Our contribution is multifaceted: (1) we develop the GEMs family using advanced architectures not previously applied systematically to Greek, such as ELECTRA, ConvBERT and ModernBERT, to counter the lack of architectural diversity; (2) we address data scarcity by constructing and releasing several large-scale, meticulously curated corpora, including a comprehensive general-domain dataset and specialized legal datasets; (3) we investigate data-centric strategies, such as the targeted repetition of high-quality legal data for domain adaptation and put forward bilingual Greek-English models (GEM) for the legal domain; and (4) we conduct a comprehensive comparative analysis on diverse downstream tasks, establishing strong new baselines for the Greek NLP community.
3 Pretraining Corpus Composition
In this research, the proposed training corpus combines legal and non-legal datasets to ensure a domain-relevant linguistically balanced representation. The legal corpora includes official government publications (ΦΕΚ)[2], the Greek part of the European Parliament Proceedings Parallel Corpus (Europarl) [22], the Greek part of European Union legislation (Eurolex) [7], the Greek Parliament Proceedings (Greekparl) [13], case law and legal interpretations (Political Reports of the Supreme Court) [1] and reference materials (Raptarchis) [31]. These sources provide a robust foundation for capturing domain-specific legal language and concepts. The emphasis on these legal corpora stems from their inherent high quality and the limited availability of alternative large-scale, curated datasets for the Modern Greek legal domain. To further enhance the model’s general language proficiency, we also incorporated nonlegal corpora, including the Greek part of Wikipedia [^1] https://dumps.wikimedia.org/elwiki/ and the Greek part of OSCAR [42], consisting of both Greek-only and Greek-English records. The statistics of the pre-training corpora are presented in Table 1.
To train our bilingual models, we created a custom 26GB bilingual corpus specifically for cross-lingual legal understanding. This unique dataset combines our proprietary Greek legal data (Table 1) with an English subset from the Pile of Law [17], resulting in a comprehensive collection of Greek and English legal, parliamentary and governmental texts.

### v2

@@ Abstract
The advancement of natural language processing for morphologically rich and moderately-resourced languages like Modern Greek has been hindered by architectural stagnation, data scarcity, and limited context processing capabilities, particularly in specialized domains such as law. In this work, we propose the Greek Embedding Models (GEMs), a new family of transformer-based language models, specifically developed to address these limitations through architectural diversity and enhanced data curation. The proposed family of models are trained on several large-scale, meticulously curated corpora, encompassing both comprehensive general-domain datasets and specialized legal collections, addressing the persistent data scarcity that has impeded Greek language modeling advancement. The proposed quality-based corpus curation methodology incorporates extensive preprocessing pipelines, sophisticated deduplication strategies and targeted repetition of high-quality legal sub-corpora to enhance domain adaptation. The GEMs family comprises both established architectures (RoBERTa and Longformer) and advanced models not previously applied to Greek (ELECTRA, ConvBERT, and ModernBERT), providing comprehensive coverage of modern transformer designs. Additionally, we introduce the first bilingual Greek-English embedding models tailored for cross-lingual legal applications. Comprehensive evaluation across three core natural language understanding benchmarks demonstrates that the proposed GEM-RoBERTa and GEM-ConvBERT achieve statistically significant performance improvements over established state-of-the-art models, with accuracy gains of up to 3.6% while conducted statistical analysis using Friedman Aligned-Ranks and Finner post-hoc tests confirms the superiority of our approach across multiple evaluation metrics.
@@ 1 Introduction
The advent of transformer-based language models vaswani2017attention ; devlin2019bert has revolutionized Natural Language Processing (NLP), establishing a dominant paradigm of large-scale pre-training followed by task-specific fine-tuning. While foundational models were initially developed for English, subsequent research has consistently shown that performance is significantly enhanced by creating monolingual models for other languages and adapting them to high-value domains, such as law (e.g., for Greek papaloukas2021multi ; apostolopoulou2022nlp and German Darji_2023 ) or biomedicine (e.g., for French labrak2023drbertrobustpretrainedmodel ). For moderately-resourced languages such as Greek, this specialization is crucial, as dedicated monolingual models like Greek-BERT koutsikakis2020greekbert have proven superior to their multilingual counterparts on a range of NLP tasks.
@@
The Greek NLP ecosystem has since evolved rapidly along several parallel tracks. Following the foundational Greek-BERT koutsikakis2020greekbert , a line of domain-specific models for the legal field emerged, including the two versions of Greek Legal BERT papaloukas2021multi ; apostolopoulou2022nlp and the architecturally superior Greek-Legal RoBERTa saketos2024greeklegalroberta , which progressively advanced the state-of-the-art on Greek legal tasks. Concurrently, separate efforts addressed this limitation for general-domain Greek with models like the Greek Longformer, which supports a 4096-token context window zaikisgreeklongformer and Meltemi-7B voukoutis2024meltemiopenlargelanguage , a large-scale generative model with an 8192-token context. However, the divergent progress, has created a fragmented landscape where the benefits of modern architectures and extended context lengths have not been systematically unified for specialized domains like Greek law.
@@
Despite this progress, further advancement in Greek NLP is hindered by several fundamental limitations. Firstly, the development of high-quality, large-scale datasets remains a significant challenge, often leading to sub-optimal model performance. Secondly, researchers also frequently neglect the inherent linguistic complexity of morphologically rich languages like Greek, where intricate syntactic structures and domain-specific terminology demand special attention. This leads to a third limitation: a lack of architectural innovation. The Greek NLP space and the legal domain in particular, has been confined to early encoder architectures like BERT and RoBERTa, often with a restrictive 512-token context length, while more efficient and powerful models such as ELECTRA clark2020electra and DeBERTa he2020deberta have not been explored. This architectural stagnation prevents the community from leveraging state-of-the-art advancements effectively.
@@
This work is motivated by the need to systematically address these challenges. We bridge the existing gaps by undertaking the first comprehensive pre-training and evaluation of a diverse suite of modern transformer architectures for the Greek language, with a specific focus on the complex legal domain as a proving ground. We specifically target this domain as its highly structured, terminologically rich and stylistically consistent language serves as a low-noise, high-signal corpus, ideal for effectively capturing the complex grammar and syntax of Greek. We propose a new family of Greek Embedding Models (GEMs), pre-trained from scratch on several meticulously curated corpora, including both a large-scale general-purpose dataset and specialized domain-specific ones. Our approach explores architectural innovations not previously applied to Greek and is the first to systematically train and evaluate a suite of Greek models with an extended 1024-token context window. Furthermore, we investigate data-centric training strategies to enhance model performance. In summary, the main contributions of this work are:
@@
  • We investigate data-centric strategies, showing that targeted repetition of high-quality legal sub-corpora is an effective strategy for domain adaptation. We also introduce the first bilingual Greek-English Embedding Models specifically tailored for the legal domain.
@@
The remainder of this work is organized as follows: Section 2 reviews related work in Greek language modeling. Section 3 presents the proposed family of model, focusing on dataset composition, corpora configurations and model architectures. Section 4 presents the results of the numerical experiments and analysis. Finally, Section 5 discusses our findings and concludes the paper.
@@ 2 Related Work
The development of effective pre-trained language models for morphologically rich languages has emerged as a fundamental challenge in advancing NLP capabilities beyond English-centric paradigms rogers2020primer ; qiu2020pre . This challenging task has been primarily driven by two complementary research directions: (i) foundational architectural advancements in transformer-based models and their optimization strategies vaswani2017attention ; devlin2019bert ; liu2019roberta and (ii) the systematic adaptation of these innovations to specific languages and specialized domains martin2020camembert ; chalkidis2020legalbert ; beltagy2019scibert . The first direction encompasses sophisticated developments in model architectures, training objectives and efficiency improvements while the second ones focuses on applying these architectural advances to non-English languages and domain-specific applications, where monolingual and specialized models consistently demonstrate superior performance over their multilingual counterparts virtanen2019multilingual ; kenton2019bert . In the remainder of this section, we provide a comprehensive review of both areas, highlighting their contributions and limitations which motivate our proposed systematic approach to Greek language model development.
@@ 2.1 Architectural Innovations in Pre-trained Language Models
The evolution of transformer-based language models has been characterized by continuous architectural refinements aimed at improving training efficiency, model capability and contextual understanding. The foundational work began with the transformer architecture vaswani2017attention and BERT’s devlin2019bert Masked Language Model (MLM) objective, which established the dominant paradigm of large-scale pre-training followed by task-specific fine-tuning. This seminal work demonstrated the effectiveness of bidirectional encoding and self-supervised learning for capturing contextual representations across diverse NLP tasks.
@@
Building upon BERT’s foundation, RoBERTa liu2019roberta introduced critical optimization improvements to the pre-training methodology, including the removal of the Next Sentence Prediction objective, dynamic masking strategies and extended training regimens. These modifications resulted in substantial performance gains across multiple benchmarks, establishing the importance of training procedure optimization alongside architectural innovation. The success of RoBERTa highlighted that careful attention to training dynamics could yield significant improvements without fundamental architectural changes.
@@
ELECTRA clark2020electra introduced a paradigm shift through its sample-efficient Replaced Token Detection (RTD) objective, which trains the model to distinguish between original and replaced tokens rather than predicting masked tokens. This approach enables the model to learn from all input positions rather than only the masked subset, resulting in improved sample efficiency and computational effectiveness. The RTD objective demonstrated particular advantages in scenarios with limited computational resources while maintaining competitive performance across downstream tasks.
@@
DeBERTa he2020deberta advanced the field through its disentangled attention mechanism, which separately encodes content and position information before combining them in attention computation. This architectural innovation, combined with enhanced mask decoder and absolute position embedding strategies, yielded substantial improvements in model performance while providing better interpretability of attention patterns. The disentangled approach proved particularly effective for tasks requiring fine-grained understanding of positional relationships and syntactic structures.
@@
The challenge of processing longer sequences was addressed by Longformer beltagy2020longformer , which introduced sparse attention patterns to handle extended contexts efficiently. By combining local windowed attention, global attention for task-specific tokens and dilated attention patterns, Longformer successfully scaled transformer models to process documents with thousands of tokens while maintaining computational tractability. This innovation proved crucial for applications requiring document-level understanding and long-range dependency modeling.
@@
ConvBERT jiang2020convbert introduced a novel hybrid approach which combines convolutional and self-attention mechanisms to improve efficiency and effectiveness. The model replaces some self-attention heads with cheaper convolutional operations, which capture local dependencies more efficiently while maintaining the global modeling capabilities of attention mechanisms. This architectural innovation reduces computational costs while often achieving superior performance, particularly on tasks requiring fine-grained local pattern recognition alongside global context understanding.
@@
Most recently, ModernBERT warner2024modernbert has emerged as a comprehensive modernization of the original BERT architecture, incorporating lessons learned from years of research into transformer optimization. ModernBERT integrates several architectural improvements including GeGLU activation functions, RMSNorm normalization, unpadding techniques for efficiency and optimized attention implementations. The model demonstrates that systematic incorporation of incremental improvements can yield substantial performance gains while maintaining compatibility with existing BERT-based infrastructure, representing a mature synthesis of architectural innovations developed over the transformer era.
@@ 2.2 Greek Language Models and Domain Specialization
The development of Greek natural language processing capabilities has followed a progressive trajectory from foundational monolingual models to specialized domain-specific applications. Koutsikakis et al. koutsikakis2020greekbert conducted the pioneering effort in modern Greek NLP by developing Greek BERT, the first monolingual BERT model specifically for the Greek language. Pre-trained on a comprehensive 29GB corpus comprising general-domain text from diverse sources, Greek BERT established crucial performance baselines and demonstrated superior results on tasks such as named entity recognition and natural language inference compared to multilingual models. This foundational work highlighted the significant benefits of language-specific pre-training for morphologically rich languages, where specialized tokenization and vocabulary development prove essential for capturing linguistic nuances.
@@
The success of Greek BERT catalyzed rapid development in domain-specific applications, particularly within the legal field. Papaloukas et al. papaloukas2021multi introduced the first Greek Legal BERT by adapting the original Greek BERT architecture through continued pre-training on a 5GB corpus of legal documents. This domain adaptation strategy demonstrated substantial improvements on legal-specific tasks, establishing the effectiveness of specialized pre-training for capturing domain-specific terminology and discourse patterns. The model’s superior performance on legal topic classification and named entity recognition tasks validated the importance of domain-focused training data curation.
@@
Apostolopoulou et al. apostolopoulou2022nlp extended this work by developing an enhanced version of Greek Legal BERT which incorporated a larger and more diverse legal corpus, including additional sources of legal documents and regulatory texts. The expanded training data provided broader coverage of legal language variations and contemporary legal discourse, resulting in improved performance metrics across multiple legal NLP tasks. This work demonstrated that systematic corpus expansion and diversification constitute effective strategies for enhancing domain-specific model capabilities.
@@
Saketos et al. saketos2024greeklegalroberta represented the next significant advancement by moving beyond the original BERT architecture and leveraging RoBERTa’s optimized pre-training methodology for Greek legal applications. Greek-Legal RoBERTa demonstrated substantial performance gains over the Greek Legal BERT variants on identical downstream legal tasks, confirming that architectural improvements and training optimizations provide complementary benefits to specialized domain adaptation. The success of this approach established that both architectural innovation and domain-specific data curation are essential components for advancing specialized NLP capabilities.
@@
Parallel developments addressed the limitation of context length processing for general-domain Greek applications. The Greek Longformer zaikisgreeklongformer introduced extended context capabilities with a 4096-token window, enabling processing of longer Greek documents while maintaining computational efficiency through sparse attention mechanisms. Similarly, Meltemi-7B voukoutis2024meltemiopenlargelanguage , a large-scale generative model, further extended context capabilities to 8192 tokens, demonstrating the feasibility of long-context modeling for Greek language applications. However, these general-domain advances remained isolated from specialized applications, creating a disconnect between architectural capabilities and domain-specific requirements.
@@ 2.3 Research Gaps and Contributions
The Greek NLP landscape is fragmented and characterized by several limitations, which impede systematic progress. A key issue is the lack of a concerted effort to create comprehensive, large-scale and high-quality datasets which are able to serve both general-purpose and specialized applications. This absence of systematically curated training corpora fundamentally limits the potential for developing robust models. In addition, the application of different model architectures has been piecemeal; while foundational models exist, more recent architectures such as ELECTRA, ModernBERT, DeBERTa and ConvBERT have not been systematically employed and evaluated for Greek language.
In this research, we address these limitations by presenting Greek Embedding Models (GEMs), a new family of models developed through a comprehensive and systematic approach. Our contribution is multifaceted: (1) we develop the GEMs family using advanced architectures, which is composed by both new-presented and advanced architectures (ELECTRA, ConvBERT and ModernBERT) as well as established (RoBERTa and LongFormer) transformer models for countering the lack of architectural diversity; (2) we address data scarcity by developing several large-scale, meticulously curated corpora, including specialized legal datasets and a comprehensive general-domain dataset; (3) finally, we develop two bilingual Greek-English Embeddings Models, which are tailed for the legal domain. The comprehensive comparative analysis on diverse downstream tasks provides empirical and statistical evidence about the effectiveness of the proposed approach, establishing strong new baselines for the Greek NLP community.
3 Development of Greek Embeddings Models
3.1 Pretraining Corpus Composition
[…]
To train our bilingual models, we created a custom 26GB bilingual corpus specifically for cross-lingual legal understanding. This unique dataset combines our proprietary Greek legal data (Table 1) with an English subset from the Pile of Law henderson2022pile , resulting in a comprehensive collection of Greek and English legal, parliamentary and governmental texts.